### v1

@@ Introduction
In recent years, multilayer Bernal-stacked graphene (AB) and graphene with rhombohedral (ABC) stacking have emerged as remarkable platforms for correlated and topological quantum phenomena [de2022cascade, li2023charge, wagner2024, henck2018flat, dong2024anomalous, bernevig2025berry]. In particular, experimental studies of rhombohedral graphene have revealed the formation of flat-band states, interaction-driven insulating phases, magnetism, and superconductivity in devices with just a few stacked layers. For example, in tetralayer or higher ABC-stacked graphene the interplay of low-density carriers, large Berry curvature, and strong Coulomb interactions has given rise to anomalous Hall and superconducting states [zhou2024layer, choi2025superconductivity, han2025signatures]. These advances underscore the power of stacking geometry, interlayer coupling, and band flattening to access exotic electronic ground states in a purely carbon-based van der Waals system.
At the same time, heterostructures coupling graphene with two-dimensional magnetic or correlated materials are rapidly expanding the playground of quantum phenomena. One prominent example is the van der Waals assembly of graphene with -RuCl3 [mashhadi2019spin, zhou2019evidence, biswas_electronic_2019, leeb2021anomalous, rizzo2020charge], a layered Mott insulating Kitaev material candidate [banerjee2016proximate, winter2017breakdown, winter2017models, trebst2022kitaev]. In these -RuCl3/graphene heretostructures, transport experiments provide evidence for both charge transfer and magnetic proximity effects [mashhadi2019spin, zhou2019evidence, leeb2021anomalous], while theoretical studies point to strain, enhanced Kitaev interactions, and interface-driven electronic reconstruction [biswas_electronic_2019, leeb2021anomalous]. The observation of nanometer-scale p–n junctions in STM measurements [balgley2022] further illustrates the richness of the interfacial physics in these systems [ojeda2025].
@@
These findings open pathways to engineer interfaces that combine flat-band physics, magnetism, and proximity effects, following a long tradition of magnetically intercalated graphite [dresselhaus_intercalation, chernyshev2022]. Bringing these threads together, the prospect of graphite intercalated with -RuCl3 (i.e., inserting layers of -RuCl3 into the van der Waals gaps of graphite or few-layer graphene) is particularly compelling. On the one hand, the flat-band and correlation physics of rhombohedral graphene suggest that modifier layers – such as intercalants – can tune the carrier density, screening, and interlayer coupling in ways that may enhance or stabilize exotic phases. On the other hand, the graphene/-RuCl3 heterostructure results show that -RuCl3 is a highly effective electron acceptor and magnetically active component with possible spin-liquid behavior in close proximity to graphene sheets. Intercalating -RuCl3 into graphite therefore provides a pathway to realizing a three-dimensional analogue of these phenomena.
@@
The synthesis of graphite intercalates can be performed using an intercalate in the vapor, liquid, or solid phase. Fabrication methods include solid-state, molten salt, two-zone vapor transport intercalation, and hydrothermal synthesis [dresselhaus_intercalation, synthesis_Hui, synthesis_matsumoto, synthesis_flandrois, synthesis_tzeng, synthesis_postnikov]. Among these, two-zone vapor transport is most widely used to intercalate volatile material into the graphite. We grew the intercalated graphite samples via a chemical vapor transport (CVT) method. Three starting constituents were sealed in a quartz ampule: 1) single crystals of graphite, in the form of natural graphite flakes from HQ graphene (99.98% purity), 2) pure -RuCl3 powder from Furuya Metals, and 3) AgCl powder from Sigma Aldrich (99.998% purity). The weight ratio used to synthesize intercalated graphite was C : -RuCl3: AgCl 0.03g : 0.2g : 0.5g. The -RuCl3 powder was pressed into a pellet to keep it together, and the AgCl powder was initially kept in an alumina crucible. We found that without excess Cl2, -RuCl3 would not intercalate into graphite. To generate a Cl2-rich atmosphere, Cl2 gas was generated by shining UV light on the AgCl powder until it decomposed, leaving behind Ag powder. We used powdered AgCl to increase the surface area exposed to UV light. Typically, the sample was irradiated with UV light for 3–5 minutes, until most of the powder turned dark brown. The AgCl section of the tube was then sealed off, leaving the graphite flakes, -RuCl3 pellet, and an enhanced Cl2 atmosphere. This is schematically shown in Fig. 2a). After the ampule was prepared, synthesis was carried out in a two-zone tube furnace. The graphite end was placed in the high temperature zone, which was held at 900 ∘C, while the -RuCl3 end was held at 830 ∘C. After 22 hours, stage 4 -RuCl3 intercalated into graphite, that we denote C-RuCl3 was formed. Continuing the reaction for a total of 48 hours resulted in stage 2 C-RuCl3. Continuing the reaction further led to mixed stage intercalation and/or disordered compounds.
@@
By utilizing the structural analogy to other transition metal trichlorides, different intercalation stackings shown in Fig. 1 are anticipated for -RuCl3-intercalated graphite [Dresselhaus2002, Cowley1956, Stumpp1977]. In Fig. 1, A, B and C denote different orientations of the graphene layers and ”” represents one -RuCl3 layer. X-ray diffraction (XRD) experiments of powdered samples resulted in two values for the c-axis and both of them show a significant expansion with respect to the c-axis of graphite (6.7 Å 25.4(2) Å and 58.2(1) Å), suggesting the successful intercalation of -RuCl3 between the layers (see Fig. S1a) and b) in the supplemental material.). From the structural analogy to other transition metal-trichlorides, we concluded the synthesized samples are stage 2 (S2) and stage 4 (S4) samples (Fig. 1b, c, and d). In the S2 sample, two graphene layers are sandwiched by two -RuCl3 layers, while in S4, four graphene layers are sandwiched by two -RuCl3 layers. XRD spots from single crystal (highly oriented polycrystalline) S2 and S4 samples can be indexed with hexagonal structures, with a small portion of residual pure graphite (see Fig. S1b) and c) in the supplemental material). The estimated a-axes are 29.62(3) Å for S2 and 29.5(A) Å for S4. The graphene and -RuCl3 unit cells alternate in the crystallographic -direction forming super lattices in both samples, where N1, N2, M1, M2 are integers.
@@
Fig. 2b) shows the temperature dependence of the longitudinal resistivity for three samples: pristine Kish graphite, Kish graphite annealed in Cl2 gas, and graphite intercalated with -RuCl3, denoted C-RuCl3. Pristine Kish graphite was used as a reference. To isolate the effects of the synthesis procedure, we also annealed graphite in a chlorine atmosphere at 900 ∘C for several days in a sealed tube without a -RuCl3 pellet ensuring any observed changes were not solely due to the annealing environment. It is important to note that the reported resistivity values should not be considered absolute, as the samples are flaky and irregular in shape. While we made careful efforts to estimate the sample dimensions, uncertainties remain. In particular, we found that using a blade to carve bar-shaped pieces sometimes alters the residual resistivity ratio [(RRR = R(300K)/R(2K)], likely due to mechanical damage.
Fig. 2c)–e) present the low-temperature magnetic field dependence of the resistivity for the Kish graphite, stage 2 and stage 4 intercalated samples. For the Kish graphite, Shubnikov–de Haas (SdH) quantum oscillations are observable at magnetic fields as low as 1.5 T. The intercalated samples exhibit SdH oscillations at T, and with much shorter periods. The frequency of the SdH oscillations was obtained by first subtracting a smooth polynomial background from the magnetoresistivity data to isolate the oscillatory component. The oscillatory part is periodic in the inverse of the magnetic field. A Fast Fourier Transform (FFT) was then performed on the oscillatory signal to extract the oscillation frequency. The frequencies we obtained for Kish graphite are T and T, with the first harmonic of also clearly observable; as shown in Fig. 2f). This is in good agreement with previous determinations of the fundamental SdH frequencies in graphite, including those obtained using other experimental techniques [Graphite_SdH_1, Graphite_SdH_2]. As shown in Fig. 2g) and h), the frequency of the SdH oscillations in the -RuCl3 intercalated graphite samples occurs around 170–176 T. The first harmonic, shown by , is also clearly observable. In addition to the main frequency, the C-RuCl3 S4 sample shows an additional SdH oscillation frequency at T.
@@
To verify this, we benchmark the strained structure against the unstrained one. We consider an unstrained C-RuCl3 structure of a 55 -RuCl3 layer intercalated into a single layer of 1212 graphene in the AA (Fig. 1a)) stacking and a strained unit cell consisting of a 22 -RuCl3 layer intercalated into a single 55 graphene layer with the same stacking type. Both unstrained and strained structures are relaxed by keeping the strong C-C bond in graphene fixed while the -RuCl3 layer is fully relaxed.
We compute the electronic properties of both type of structures within DFT and compare them with those of pristine single-layer graphene and bulk -RuCl3. By aligning the electronic bands of the unstrained C-RuCl3 with those of the pristine materials, we observe no significant hybridization between the graphene and -RuCl3 layers; instead, mostly charge transfer takes place (see Supplementary Material), in agreement with previous studies on similar heterostructures [biswas_electronic_2019, leeb2021anomalous, rossi2023direct, razpopov2024ab]. Because the hybridization is effectively very weak, the applied strain in the simplified unit cell impacts only the -RuCl3 electronic structure, leaving the graphene bands unchanged except for a band shift due to charge transfer (see below). Using the graphene--RuCl3 separation suggested by X-ray data, we find nearly identical charge depletion in both, strained and unstrained systems. We also tested different relative orientations and found that they produce no significant changes in the charge transfer (see Supplementary Material).
@@
Following this approach, we model the experimental structures determined by XRD by constructing four geometries containing a tensile-strained 22 -RuCl3 layer intercalated with a 55 graphene supercell in the stacking arrangements AA, ABAB, ABABABAB and ABCAABCA as shown in Fig. 1. The theoretical structures ABAB correspond to the S2 sample, and ABABABAB and ABCAABCA to the S4 sample. Here, to reduce computational expense, we replicate the graphene stacking sequence after each -RuCl3 layer, simplifying the original stacking from AABB to ABAB, and ABBCABBC (ABBABAAB) to ABCAABCA (ABABABAB). This simplification is justified because the charge transfer is insensitive to the relative orientation between -RuCl3 and graphene, and the graphene sections are well separated by the intervening -RuCl3 layer.
Our Bader analysis of the charge redistribution (Fig. 3) shows that the graphene sheets donate charge to the -RuCl3 sheet, and the charge mostly accumulates at the interface (Fig. 3 a)). Furthermore, the total charge accumulated on the -RuCl3 layer increases with the number of intercalated graphene layers (see Supplementary Material for details). A similar trend has recently been reported in multilayer graphene heterostructures [falorsi2025interlayer]. The magnitude of the charge accumulation is consistent with previous studies of related -RuCl3/graphene systems [biswas_electronic_2019, balgley2022, razpopov2024ab]. When comparing the charge transfer between the different ABABABAB and ABCAABCA stacking configurations, we find no significant differences.
In Fig. 3b), we present the planar-averaged charge density along the c direction. In the AA intercalation, every graphene layer exhibits the same charge depletion because of the periodicity of the unit cell. This also holds for the ABAB intercalation, as each layer faces a -RuCl3 surface. In contrast, for the ABABABAB or ABCAABCA configurations, charge depletion is concentrated primarily on the interface graphene layers, while the inner graphene layers remain screened.
@@
To understand the origin of the measured quantum oscillations, we estimate the corresponding frequencies using the Onsager relation [onsager1952interpretation], which links the frequencies to the extremal Fermi surface cross-sectional area . These are compared with the experimental data in Fig. 2g) and h). Since the magnetic field in the experiments is applied along the crystal -axis, we consider Fermi surface cross sections perpendicular to . Electronic structures and corresponding Fermi surfaces are calculated for each intercalated crystal.
@@
The charge redistribution at the interface between -RuCl3 and the adjacent graphene layer is mostly localized at the Cl atoms as shown in Fig. 3a) with -RuCl3 remaining rather insulating [zheng2024incommensurate]. Therefore, only the graphene layers are expected to contribute to the observed quantum oscillations. We compute the graphene layers electronic properties directly from the intercalated structures using non-spin-polarized GGA. The band structures are unfolded and projected onto graphene states, then compared to pristine, undoped graphene (Fig. 4). Note the formation of a flat band with the number of stacked graphene layers, as has been observed in rombohedral graphene stacking [aoki2007dependence, johansson2014multiple, yagi2018low, henck2018flat, zhang2025layer]. In all intercalated cases, charge transfer induces hole pockets around the K points.
@@
In the AA (Fig. 4a)) and ABAB(Fig. 4b)) intercalations, the pristine graphene bands are preserved with a global Fermi energy shift of 0.68 eV and 0.48 eV, respectively, due to equivalent hole doping of the graphene layers (Fig. 3b)). The smaller shift in the bilayer ABAB case arises because the transferred charge is distributed across two bands, whereas in the single-layer AA case only one band contributes to the depletion. In the four-layer ABABABAB and ABCAABCA stackings, the energy shift is non-uniform, reflecting the layer-dependent charge redistribution (Fig. 4c), d)). The pristine four-layer structures show four energy-separated bands, in agreement with previous studies [aoki2007dependence, johansson2014multiple, yagi2018low, henck2018flat, zhang2025layer], while in the intercalated systems, two of these bands overlap or nearly overlap due to doping. The band geometry around K also depends on the stacking, directly influencing the Fermi surface.
@@
The Fermi surfaces are computed by unfolding the bands and projecting onto graphene states, focusing on the region around K where the bands cross the Fermi level (Fig. 5). The AA intercalation yields a single hole pocket, consistent with prior experimental observations [mashhadi2019spin]. ABAB produces two pockets of different sizes, while ABABABAB and ABCAABCA exhibit three hole pockets. In the ABABABAB, the three pockets are distinct, with the two outer bands overlapping, forming a larger pocket plus a smaller one. In the ABCAABCA, the three pockets are distinct, with the two outer bands nearly overlapping. Thus, the stacking sequence controls the number and size of Fermi hole pockets. We compute their areas and the corresponding quantum oscillation frequencies using the Onsager equation, and compare the calculated quantum oscillation frequencies with the experimentally measured ones ( T, T).

### v2

@@ Introduction
In recent years, multilayer Bernal-stacked graphene (AB) and graphene with rhombohedral (ABC) stacking have emerged as remarkable platforms for correlated and topological quantum phenomena [1, 2, 3, 4, 5, 6]. In particular, experimental studies of rhombohedral graphene have revealed the formation of flat-band states, interaction-driven insulating phases, magnetism, and superconductivity in devices with just a few stacked layers. For example, in tetralayer or higher ABC-stacked graphene the interplay of low-density carriers, large Berry curvature, and strong Coulomb interactions has given rise to anomalous Hall and superconducting states [7, 8, 9]. These advances underscore the power of stacking geometry, interlayer coupling, and band flattening to access exotic electronic ground states in a purely carbon-based van der Waals system.
At the same time, heterostructures coupling graphene with two-dimensional magnetic or correlated materials are rapidly expanding the playground of quantum phenomena. One prominent example is the van der Waals assembly of graphene with -RuCl3 [10, 11, 12, 13, 14], a layered Mott insulating Kitaev material candidate [15, 16, 17, 18]. In these -RuCl3/graphene heretostructures, transport experiments provide evidence for both charge transfer and magnetic proximity effects [10, 11, 13], while theoretical studies point to strain, enhanced Kitaev interactions, and interface-driven electronic reconstruction [12, 13]. The observation of nanometer-scale p–n junctions in STM measurements [19] further illustrates the richness of the interfacial physics in these systems [20].
@@
These findings open pathways to engineer interfaces that combine flat-band physics, magnetism, and proximity effects, following a long tradition of magnetically intercalated graphite [21, 22]. Bringing these threads together, the prospect of graphite intercalated with -RuCl3 (i.e., inserting layers of -RuCl3 into the van der Waals gaps of graphite or few-layer graphene) is particularly compelling. On the one hand, the flat-band and correlation physics of rhombohedral graphene suggest that modifier layers – such as intercalants – can tune the carrier density, screening, and interlayer coupling in ways that may enhance or stabilize exotic phases. On the other hand, the graphene/-RuCl3 heterostructure results show that -RuCl3 is a highly effective electron acceptor and magnetically active component with possible spin-liquid behavior in close proximity to graphene sheets. Intercalating -RuCl3 into graphite therefore provides a pathway to realizing a three-dimensional analogue of these phenomena, as illustrated in Figure 1.
@@
The synthesis of graphite intercalates can be performed using an intercalate in the vapor, liquid, or solid phase. Fabrication methods include solid-state, molten salt, two-zone vapor transport intercalation, and hydrothermal synthesis [21, 23, 24, 25, 26, 27]. Among these, two-zone vapor transport is most widely used to intercalate volatile material into the graphite. We grew the intercalated graphite samples via a chemical vapor transport (CVT) method. Three starting constituents were sealed in a quartz ampule: 1) single crystals of graphite, in the form of natural graphite flakes from HQ graphene (99.98% purity), 2) pure -RuCl3 powder from Furuya Metals, and 3) AgCl powder from Sigma Aldrich (99.998% purity). The weight ratio used to synthesize intercalated graphite was C : -RuCl3: AgCl 0.03g : 0.2g : 0.5g. The -RuCl3 powder was pressed into a pellet to keep it together, and the AgCl powder was initially kept in an alumina crucible. We found that without excess Cl2, -RuCl3 would not intercalate into graphite. To generate a Cl2-rich atmosphere, Cl2 gas was generated by shining UV light on the AgCl powder until it decomposed, leaving behind Ag powder. We used powdered AgCl to increase the surface area exposed to UV light. Typically, the sample was irradiated with UV light for 3–5 minutes, until most of the powder turned dark brown. The AgCl section of the tube was then sealed off, leaving the graphite flakes, -RuCl3 pellet, and an enhanced Cl2 atmosphere. This is schematically shown in Figure 2a. After the ampule was prepared, synthesis was carried out in a two-zone tube furnace. The graphite end was placed in the high temperature zone, which was held at 900 ∘C, while the -RuCl3 end was held at 830 ∘C. After 22 hours, stage 4 -RuCl3 intercalated into graphite, that we denote C-RuCl3 was formed. Continuing the reaction for a total of 48 hours resulted in stage 2 C-RuCl3. Continuing the reaction further led to mixed stage intercalation and/or disordered compounds.
@@
By utilizing the structural analogy to other transition metal trichlorides, different intercalation stackings shown in Figure 1 are anticipated for -RuCl3-intercalated graphite [28, 29, 30]. In Figure 1, A, B and C denote different orientations of the graphene layers and ”” represents one -RuCl3 layer. X-ray diffraction (XRD) experiments of powdered samples resulted in two values for the c-axis and both of them show a significant expansion with respect to the c-axis of graphite (6.7 Å 25.4(2) Å and 58.2(1) Å), suggesting the successful intercalation of -RuCl3 between the layers (see Figures S1a and b in the supplemental material.). From the structural analogy to other transition metal-trichlorides, we concluded the synthesized samples are stage 2 (S2) and stage 4 (S4) samples (Figure 1b, c, and d). In the S2 sample, two graphene layers are sandwiched by two -RuCl3 layers, while in S4, four graphene layers are sandwiched by two -RuCl3 layers. XRD spots from single crystal (highly oriented polycrystalline) S2 and S4 samples can be indexed with hexagonal structures, with a small portion of residual pure graphite (see Figures S1b and c in the supplemental material). The estimated a-axes are 29.62(3) Å for S2 and 29.5(A) Å for S4. The graphene and -RuCl3 unit cells alternate in the crystallographic -direction forming super lattices in both samples, where N1, N2, M1, M2 are integers.
@@
Figure 2b shows the temperature dependence of the longitudinal resistivity for three samples: pristine Kish graphite, Kish graphite annealed in Cl2 gas, and graphite intercalated with -RuCl3, denoted C-RuCl3. Pristine Kish graphite was used as a reference. To isolate the effects of the synthesis procedure, we also annealed graphite in a chlorine atmosphere at 900 ∘C for several days in a sealed tube without a -RuCl3 pellet ensuring any observed changes were not solely due to the annealing environment. It is important to note that the reported resistivity values should not be considered absolute, as the samples are flaky and irregular in shape. While we made careful efforts to estimate the sample dimensions, uncertainties remain. In particular, we found that using a blade to carve bar-shaped pieces sometimes alters the residual resistivity ratio [(RRR = R(300K)/R(2K)], likely due to mechanical damage.
Figure 2c–e present the low-temperature magnetic field dependence of the resistivity for the Kish graphite, stage 2 and stage 4 intercalated samples. For the Kish graphite, Shubnikov–de Haas (SdH) quantum oscillations are observable at magnetic fields as low as 1.5 T. The intercalated samples exhibit SdH oscillations at T, and with much shorter periods. The frequency of the SdH oscillations was obtained by first subtracting a smooth polynomial background from the magnetoresistivity data to isolate the oscillatory component. The oscillatory part is periodic in the inverse of the magnetic field. A Fast Fourier Transform (FFT) was then performed on the oscillatory signal to extract the oscillation frequency. The frequencies we obtained for Kish graphite are T and T, with the first harmonic of also clearly observable; as shown in Figure 2f. This is in good agreement with previous determinations of the fundamental SdH frequencies in graphite, including those obtained using other experimental techniques [31, 32]. As shown in Figure 2g and h, the frequency of the SdH oscillations in the -RuCl3 intercalated graphite samples occurs around 170–176 T. The first harmonic, shown by , is also clearly observable. In addition to the main frequency, the C-RuCl3 S4 sample shows an additional SdH oscillation frequency at T.
@@
To verify this, we benchmark the strained structure against the unstrained one. We consider an unstrained C-RuCl3 structure of a 55 -RuCl3 layer intercalated into a single layer of 1212 graphene in the AA (Figure 1a) stacking and a strained unit cell consisting of a 22 -RuCl3 layer intercalated into a single 55 graphene layer with the same stacking type. Both unstrained and strained structures are relaxed by keeping the strong C-C bond in graphene fixed while the -RuCl3 layer is fully relaxed.
We compute the electronic properties of both type of structures within DFT and compare them with those of pristine single-layer graphene and bulk -RuCl3. By aligning the electronic bands of the unstrained C-RuCl3 with those of the pristine materials, we observe no significant hybridization between the graphene and -RuCl3 layers; instead, mostly charge transfer takes place (see Supplementary Material), in agreement with previous studies on similar heterostructures [12, 13, 34, 35]. Because the hybridization is effectively very weak, the applied strain in the simplified unit cell impacts only the -RuCl3 electronic structure, leaving the graphene bands unchanged except for a band shift due to charge transfer (see below). Using the graphene--RuCl3 separation suggested by X-ray data, we find nearly identical charge depletion in both, strained and unstrained systems. We also tested different relative orientations and found that they produce no significant changes in the charge transfer (see Supplementary Material).
@@
Following this approach, we model the experimental structures determined by XRD by constructing four geometries containing a tensile-strained 22 -RuCl3 layer intercalated with a 55 graphene supercell in the stacking arrangements AA, ABAB, ABABABAB and ABCAABCA as shown in Figure 1. The theoretical structures ABAB correspond to the S2 sample, and ABABABAB and ABCAABCA to the S4 sample. Here, to reduce computational expense, we replicate the graphene stacking sequence after each -RuCl3 layer, simplifying the original stacking from AABB to ABAB, and ABBCABBC (ABBABAAB) to ABCAABCA (ABABABAB). This simplification is justified because the charge transfer is insensitive to the relative orientation between -RuCl3 and graphene, and the graphene sections are well separated by the intervening -RuCl3 layer.
Our Bader analysis of the charge redistribution (Figure 3) shows that the graphene sheets donate charge to the -RuCl3 sheet, and the charge mostly accumulates at the interface (Figure 3a). Furthermore, the total charge accumulated on the -RuCl3 layer increases with the number of intercalated graphene layers (see Supplementary Material for details). A similar trend has recently been reported in multilayer graphene heterostructures [36]. The magnitude of the charge accumulation is consistent with previous studies of related -RuCl3/graphene systems [12, 19, 35]. When comparing the charge transfer between the different ABABABAB and ABCAABCA stacking configurations, we find no significant differences.
In Figure 3b, we present the planar-averaged charge density along the c direction. In the AA intercalation, every graphene layer exhibits the same charge depletion because of the periodicity of the unit cell. This also holds for the ABAB intercalation, as each layer faces a -RuCl3 surface. In contrast, for the ABABABAB or ABCAABCA configurations, charge depletion is concentrated primarily on the interface graphene layers, while the inner graphene layers remain screened.
@@
To understand the origin of the measured quantum oscillations, we estimate the corresponding frequencies using the Onsager relation [37], which links the frequencies to the extremal Fermi surface cross-sectional area . These are compared with the experimental data in Figure 2g and h. Since the magnetic field in the experiments is applied along the crystal -axis, we consider Fermi surface cross sections perpendicular to . Electronic structures and corresponding Fermi surfaces are calculated for each intercalated crystal.
@@
The charge redistribution at the interface between -RuCl3 and the adjacent graphene layer is mostly localized at the Cl atoms as shown in Figure 3a with -RuCl3 remaining rather insulating [38]. Therefore, only the graphene layers are expected to contribute to the observed quantum oscillations. We compute the graphene layers electronic properties directly from the intercalated structures using non-spin-polarized GGA. The band structures are unfolded and projected onto graphene states, then compared to pristine, undoped graphene (Figure 4). Note the formation of a flat band with the number of stacked graphene layers, as has been observed in rombohedral graphene stacking [39, 40, 41, 4, 42]. In all intercalated cases, charge transfer induces hole pockets around the K points.
@@
In the AA (Figure 4a) and ABAB(Figure 4b) intercalations, the pristine graphene bands are preserved with a global Fermi energy shift of 0.68 eV and 0.48 eV, respectively, due to equivalent hole doping of the graphene layers (Figure 3b). The smaller shift in the bilayer ABAB case arises because the transferred charge is distributed across two bands, whereas in the single-layer AA case only one band contributes to the depletion. In the four-layer ABABABAB and ABCAABCA stackings, the energy shift is non-uniform, reflecting the layer-dependent charge redistribution (Figure 4c, d). The pristine four-layer structures show four energy-separated bands, in agreement with previous studies [39, 40, 41, 4, 42], while in the intercalated systems, two of these bands overlap or nearly overlap due to doping. The band geometry around K also depends on the stacking, directly influencing the Fermi surface.
@@
The Fermi surfaces are computed by unfolding the bands and projecting onto graphene states, focusing on the region around K where the bands cross the Fermi level (Figure 5). The AA intercalation yields a single hole pocket, consistent with prior experimental observations [10]. ABAB produces two pockets of different sizes, while ABABABAB and ABCAABCA exhibit three hole pockets. In the ABABABAB, the three pockets are distinct, with the two outer bands overlapping, forming a larger pocket plus a smaller one. In the ABCAABCA, the three pockets are distinct, with the two outer bands nearly overlapping. Thus, the stacking sequence controls the number and size of Fermi hole pockets. We compute their areas and the corresponding quantum oscillation frequencies using the Onsager equation, and compare the calculated quantum oscillation frequencies with the experimentally measured ones ( T, T).